A string in a double inclined plane connects two bodies. If the 6.0 π‘˜π‘” mass starts from the top of the plane and the length is 8 m. Find the velocity of the 6 π‘˜π‘” body when it reaches the bottom of the plane.Β 


1
Expert's answer
2021-10-03T13:22:09-0400

Assume the mass of the second body is M>m=6 kg. Assume the plane with the 6-kg body is vertical and the plane on which the other body rests is ΞΈ above the horizontal.

Applying Newton's second law for both bodies, we have


"-ma=T-mg,\\
Ma=T-Mg\sin\theta.\\\space\\
Ma=mg-ma-Mg\sin\theta,\\
a(M+m)=g(m-M\sin\theta),\\\space\\
a=\frac{g(m-M\sin\theta)}{M+m}."

The velocity near the ground is


"v=\sqrt{2gh\frac{(m-M\sin\theta)}{(M+m)}}."